# TDSM: Triplet Diffusion for Skeleton-Text Matching in Zero-Shot Action Recognition

### v1

@@ Abstract
We firstly present a diffusion-based action recognition with zero-shot learning for skeleton inputs. In zero-shot skeleton-based action recognition, aligning skeleton features with the text features of action labels is essential for accurately predicting unseen actions. Previous methods focus on direct alignment between skeleton and text latent spaces, but the modality gaps between these spaces hinder robust generalization learning. Motivated from the remarkable performance of text-to-image diffusion models, we leverage their alignment capabilities between different modalities mostly by focusing on the training process during reverse diffusion rather than using their generative power. Based on this, our framework is designed as a Triplet Diffusion for Skeleton-Text Matching (TDSM) method which aligns skeleton features with text prompts through reverse diffusion, embedding the prompts into the unified skeleton-text latent space to achieve robust matching. To enhance discriminative power, we introduce a novel triplet diffusion (TD) loss that encourages our TDSM to correct skeleton-text matches while pushing apart incorrect ones. Our TDSM significantly outperforms the very recent state-of-the-art methods with large margins of 2.36%-point to 13.05%-point, demonstrating superior accuracy and scalability in zero-shot settings through effective skeleton-text matching.
@@ 1 Introduction
Figure 1: Motivation of the proposed TDSM pipeline. Previous methods rely on direct alignment between skeleton and text latent spaces, but suffer from modality gaps that limit generalization. Our TDSM overcomes this challenge by leveraging a reverse diffusion process to embed text prompts into the unified skeleton-text latent space, ensuring more effective cross-modal alignment.
Despite these benefits, scaling skeleton-based recognition to large action sets remains challenging. Traditional methods [63, 8, 12, 70, 11, 67, 6, 9] require training on annotated data for each action, which is impractical for a large number of possible actions. Zero-shot skeleton-based action recognition [47, 16, 69, 71, 32, 7, 61, 26, 30] addresses this by enabling predictions for unseen actions without requiring explicit training data, making it valuable for applications such as surveillance, robotics, and human-computer interaction, where continuous learning is infeasible [57, 15].
@@
Zero-shot models offer the flexibility for unseen data by aligning pre-learned skeleton features with text-based descriptions corresponding to action labels, ensuring their scalability in real-world scenarios. However, achieving the effective alignment between skeleton data and text features entails significant challenges. While skeleton data captures temporal and spatial motion patterns, the text descriptions for action labels carry high-level semantic information. This modality gap makes it difficult to align their corresponding latent spaces effectively, hindering the generalization learning for unseen actions—particularly when distinguishing between semantically similar actions (e.g., “Throw” vs. “Shoot”).
Fig. 1 illustrates the key differences between previous methods and our proposed method. The previous methods [47, 16, 69, 71, 32, 7, 61, 26, 30] aim at directly aligning the different features of skeletons and text descriptions for action labels between the skeleton and text latent spaces, which struggles to generalize due to the inherent differences between these modalities in different latent spaces. On the other hand, our TDSM leverages a reverse diffusion training scheme to merge the skeleton features with their corresponding text prompts, which leads to discriminatively fused features in a unified latent space.
@@
Inspired by text-to-image diffusion models [45, 14], we propose a novel framework: Triplet Diffusion for Skeleton-Prompt Matching (TDSM). Unlike to be a generative model, our TDSM leverages the reverse diffusion training scheme to merge skeleton features with text prompts which aims at achieving merged discriminative (not generative) features for improved action recognition tasks. More specifically, our TDSM learns to denoise noisy skeleton features conditioned on the corresponding text prompts, embedding the prompts into a merged skeleton-text latent space to ensure effective implicit alignment. This approach avoids the challenges of direct latent space alignment while enhancing robustness. Additionally, we introduce a novel triplet diffusion (TD) loss, which encourages tighter alignment for correct skeleton-text pairs and pushes apart incorrect ones, further improving the model’s discriminative power. As an additional benefit, the diffusion process, driven by the random noise added during training, acts as a natural regularization mechanism. This prevents overfitting and enhances the model’s ability to generalize effectively to unseen actions. Our contributions are threefold:
@@
  • We firstly present a diffusion-based action recognition with zero-shot learning for skeleton inputs, called a Triplet Diffusion for Skeleton-Text Matching (TDSM) which is the first framework to apply diffusion models and to implicitly align the skeleton features with text prompts by fully taking the advantage of excellent text-image correspondence learning in generative diffusion process, thus being able to learn fused discriminative features in a unified latent space.
@@
  • We propose a novel triplet diffusion (TD) loss to enhance the model’s discriminative power by ensuring accurate denoising for correct skeleton-text pairs while suppressing incorrect pairs.
@@ 2 Related Work
2.1 Skeleton-based Action Recognition
Traditional skeleton-based action recognition assumes fully annotated training and test datasets, in contrast to other skeleton-based action recognition methods under zero-shot settings which aim to recognize unseen classes without explicit training samples. Early methods [66, 34, 72, 31] employed RNN-based models to capture the temporal dynamics of skeleton sequences. Subsequent studies [3, 62, 23, 12] explored CNN-based approaches, transforming skeleton data into pseudo-images. Recent advancements leverage graph convolutional networks (GCNs) [6, 67, 28, 9, 60, 36, 27, 70] to effectively represent the graph structures of skeletons, comprising joints and bones. ST-GCN [63] introduced graph convolutions along the skeletal axis combined with 1D temporal convolutions to capture motion over time. Shift-GCN [8] improved computational efficiency by implementing shift graph convolutions. Building on these methods, transformer-based models [68, 39, 13, 55, 11] have been proposed to address the limited receptive field of GCNs by capturing global skeletal-temporal dependencies. In this work, we adopt ST-GCN [63] and Shift-GCN [8] to extract skeletal-temporal representations from skeleton data, transforming input skeleton sequences into a latent space for further processing in the proposed framework.
@@ 2.2 Zero-shot Skeleton-based Action Recognition
Most of the existing works focus on aligning the skeleton latent space with the text latent space. These approaches can be categorized broadly into VAE-based methods [47, 16, 30, 32] and contrastive learning-based methods [69, 71, 7, 26, 61].
@@
While most existing methods rely on direct alignment between skeleton and text latent spaces, they often struggle with generalization due to inherent differences between the two modalities. To address these limitations, our TDSM leverages a novel triplet diffusion approach. It uses text prompts to guide the reverse diffusion process, embedding these prompts into the unified skeleton-text latent space for more effective implicit alignment via feature fusion. Also, our triplet diffusion loss, inspired by the triplet loss [19], further enhances the model’s discriminative power by promoting correct alignments and penalizing incorrect ones.
@@ 2.3 Diffusion Models
Diffusion models have become a fundamental to generative tasks by learning to reverse a noise-adding process for original data recovery. Denoising Diffusion Probabilistic Models (DDPMs) [18] introduced a step-by-step denoising framework, enabling the modeling of complex data distributions and establishing the foundation for diffusion-based generative models. Building on this, Latent Diffusion Models (LDMs) [45, 14] improve computational efficiency by operating in lower-dimensional latent spaces while maintaining high-quality outputs. LDMs have been particularly successful in text-to-image (T2I) generation, showing the potential of diffusion models for cross-modal alignment tasks. In the reverse diffusion process, LDMs employ a denoising U-Net [46] where text prompts are integrated with image features through cross-attention blocks, effectively guiding the model to align the two modalities. Further extending this line of research, Diffusion Transformers (DiTs) [41] integrate transformer architectures into diffusion processes, capturing long-range dependencies across spatial dimension. In this work, we leverage the aligned fusion capabilities of diffusion models, focusing on the learning process during reverse diffusion rather than their generative power. Specifically, we utilize a DiT-based network as a denoising model, where text prompts guide the denoising of noisy skeleton features. This approach embeds text prompts into the unified latent space in the reverse diffusion process, ensuring robust fusion of the two modalities and enabling effective generalization to unseen actions in zero-shot recognition settings.
@@ 3.1 Overview of TDSM
$\mathcal{D}_{\text{train}}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N},\quad y_{i}\in\mathcal{Y},$ (1)
@@
$\mathcal{D}_{\text{test}}=\{(\mathbf{X}_{j}^{u},y_{j}^{u})\}_{j=1}^{N_{u}},\quad y_{j}^{u}\in\mathcal{Y}_{u},$ (2)
@@
Fig. 2 provides an overview of our training framework of TDSM. The framework consists of three main components: (i) skeleton and text encoders, (ii) the forward process, and (iii) the reverse process. As detailed in Sec. 3.2, the pretrained skeleton encoder $\mathcal{E}_{x}$ and text encoder $\mathcal{E}_{d}$ embed the skeleton inputs $\mathbf{X}$ and prompt input $\mathbf{d}$ with an action label $y$ into their respective feature spaces, producing the skeleton feature $\mathbf{z}_{x}$ and two types of text features: the global text feature $\mathbf{z}_{g}$ and the local text feature $\mathbf{z}_{l}$. The skeleton feature $\mathbf{z}_{x}$ undergoes the forward process, where noise $\bm{\epsilon}$ is added to it. In the reverse process, as described in Sec. 3.3, the Diffusion Transformer $\mathcal{T}_{\text{diff}}$ predicts the noise $\hat{\bm{\epsilon}}$. The training objective function that ensures the TDSM to learn robust discriminative power is discussed in Sec. 3.3. Finally, Sec. 3.4 explains the strategy used during the inference phase to predict the correct label ${\hat{y}}^{u}$ for unseen actions.
@@ 3.2 Embedding Skeleton and Prompt Input
Following the LDMs [45, 14], we perform the diffusion process in a compact latent space by projecting both skeleton data and prompt into their respective feature spaces. For skeleton data, we adopt GCNs, specifically ST-GCN [63] and Shift-GCN [8], as the architecture for the skeleton encoder $\mathcal{E}_{x}$ that is trained on $\mathcal{D}_{\text{train}}$ using a cross-entropy loss:
@@
For text encoder $\mathcal{E}_{d}$, we leverage the text prompts provided by SA-DVAE [32] to capture rich semantic information about the action labels. Each ground truth (GT) label is associated with a prompt $\mathbf{d}_{p}$, while a wrong label (negative sample) is assigned a prompt $\mathbf{d}_{n}$. To encode these prompts, we utilize a pretrained text encoder, such as CLIP [44, 21], which provides two types of output features: a global text feature $\mathbf{z}_{g}$ and a local text feature $\mathbf{z}_{l}$. The text encoder’s output for a given prompt d can be represented as:
@@
For each GT label (positive sample), the text encoder $\mathcal{E}_{d}$ extracts both the global and local text features, denoted as $\mathbf{z}_{g,p}$ and $\mathbf{z}_{l,p}$, respectively. Similarly, for each wrong label (negative sample), the encoder extracts the features $\mathbf{z}_{g,n}$ and $\mathbf{z}_{l,n}$. These four features later guide the diffusion process by conditioning the denoising of noisy skeleton features.
@@ 3.3 Diffusion Process
$\mathbf{z}_{x,t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{x}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},$ (6)
where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ controls the noise level at step $t$ [18].
@@
$\hat{\bm{\epsilon}}=\mathcal{T}_{\text{diff}}\left(\mathbf{z}_{x,t},\mathbf{z}_{g},\mathbf{z}_{l},t\right).$ (7)
Using the shared weights in $\mathcal{T}_{\text{diff}}$, we predict $\hat{\bm{\epsilon}}_{p}$ for positive features $\left(\mathbf{z}_{g,p},\mathbf{z}_{l,p}\right)$ and $\hat{\bm{\epsilon}}_{n}$ for negative features $\left(\mathbf{z}_{g,n},\mathbf{z}_{l,n}\right)$.
Diffusion Transformer $\mathcal{T}_{\text{diff}}$. The Diffusion Transformer $\mathcal{T}_{\text{diff}}$ takes $\mathbf{z}_{x,t}$, $\mathbf{z}_{g}$, $\mathbf{z}_{l}$, and $t$ as inputs (Fig. 2). These inputs are embedded into corresponding feature representations $\mathbf{f}_{x,t}$, $\mathbf{f}_{c}$, and $\mathbf{f}_{l}$ as follows:
$\begin{split}&\mathbf{f}_{x,t}=\mathsf{Linear}(\mathbf{z}_{x,t})+\mathsf{PE}_{x},\\ &\mathbf{f}_{c}=\mathsf{Linear}(\mathsf{TE}_{t})+\mathsf{Linear}(\mathbf{z}_{g}),\\ &\mathbf{f}_{l}=\mathsf{Linear}(\mathbf{z}_{l})+\mathsf{PE}_{l},\\ \end{split}$ (8)
where $\mathsf{PE}_{x}$ and $\mathsf{PE}_{l}$ are positional embeddings applied to the feature maps, capturing spatial positional information, while $\mathsf{TE}_{t}$ is a timestep embedding [53] that maps the scalar $t$ to a higher-dimensional space. The embedded features $\mathbf{f}_{x,t}$, $\mathbf{f}_{c}$, and $\mathbf{f}_{l}$ are then passed through $B$ CrossDiT Blocks, followed by a Layer Normalization ($\mathsf{LN}$) and a final $\mathsf{Linear}$ layer to predict the noise $\hat{\bm{\epsilon}}\in\mathbb{R}^{M_{x}\times C}$.
CrossDiT Block. The CrossDiT Block facilitates interaction between skeleton and text features, enhancing fusion through effective feature modulation [58, 42] and multi-head self-attention [54]. Fig. 3 shows a detail structure of our CrossDiT Block. Built upon the DiTs architecture [41, 14], it leverages modulation techniques and self-attention mechanisms to efficiently capture the dependencies across these modalities. The skeleton feature $\mathbf{f}_{x}$ and local text feature $\mathbf{f}_{l}$ are first modulated separately using $\mathsf{Scale}$-$\mathsf{Shift}$ and $\mathsf{Scale}$ operations as:
$\begin{split}&\left[\bm{\alpha}_{x}\mid\bm{\beta}_{x}\mid\bm{\gamma}_{x}\>|\>\bm{\alpha}_{l}\mid\bm{\beta}_{l}\mid\bm{\gamma}_{l}\right]=\mathsf{Linear}(\mathbf{f}_{c}),\\ &\mathsf{Scale}\text{-}\mathsf{Shift}:\;\mathbf{f}_{i}\leftarrow(1+\bm{\gamma}_{i})\odot\mathbf{f}_{i}+\bm{\beta}_{i},\\ &\mathsf{Scale}:\;\mathbf{f}_{i}\leftarrow\bm{\alpha}_{i}\odot\mathbf{f}_{i},\\ \end{split}$ (9)
where $i\in\left\{x,l\right\}$ denotes the skeleton or local text feature, respectively. The parameters $\bm{\alpha}$, $\bm{\beta}$, and $\bm{\gamma}$ are conditioned on the global text feature $\mathbf{z}_{g}$ and timestep $t$, allowing the block to modulate feature representations effectively. Also, we compute query, key, and value matrices for both skeleton and local text features separately:
$[\mathbf{q}_{i}\mid\mathbf{k}_{i}\mid\mathbf{v}_{i}]=\mathsf{Linear}(\mathbf{f}_{i}).$ (10)
These matrices are token-wise concatenated and fed into a multi-head self-attention module, followed by a split to retain token-specific information as:
$[\mathbf{f}_{x}\mid\mathbf{f}_{l}]\leftarrow\mathsf{SoftMax}\left(\left[\mathbf{q}_{x}\mid\mathbf{q}_{l}\right]\left[\mathbf{k}_{x}\mid\mathbf{k}_{l}\right]^{\mathsf{T}}\right)\left[\mathbf{v}_{x}\mid\mathbf{v}_{l}\right].$ (11)
By leveraging the attention from skeleton, timestep, and text features, the CrossDiT Block ensures efficient interaction between modalities, promoting the skeleton-text fusion for discriminative feature learning and improved generalization to unseen actions.
Figure 3: A detail structure of our CrossDiT Block.
Loss function design. The overall training objective combines a diffusion loss and our triplet diffusion (TD) loss to promote both effective noise prediction and discriminative alignment (fusion). The total loss is defined as:
@@
where $\mathcal{L}_{\text{diff}}$ ensures accurate denoising, and $\mathcal{L}_{\text{TD}}$ enhances the ability to differentiate between correct and incorrect label predictions. The diffusion loss $\mathcal{L}_{\text{diff}}$ [18] is given by:
@@
Figure 4: Inference framework of our TDSM for zero-shot skeleton-based action recognition.

### v3

@@ Abstract
In zero-shot skeleton-based action recognition (ZSAR), aligning skeleton features with the text features of action labels is essential for accurately predicting unseen actions. ZSAR faces a fundamental challenge in bridging the modality gap between the two-kind features, which severely limits generalization to unseen actions. Previous methods focus on direct alignment between skeleton and text latent spaces, but the modality gaps between these spaces hinder robust generalization learning. Motivated by the success of diffusion models in multi-modal alignment (e.g., text-to-image, text-to-video), we firstly present a diffusion-based skeleton-text alignment framework for ZSAR. Our approach, Triplet Diffusion for Skeleton-Text Matching (TDSM), focuses on cross-alignment power of diffusion models rather than their generative capability. Specifically, TDSM aligns skeleton features with text prompts by incorporating text features into the reverse diffusion process, where skeleton features are denoised under text guidance, forming a unified skeleton-text latent space for robust matching. To enhance discriminative power, we introduce a triplet diffusion (TD) loss that encourages our TDSM to correct skeleton-text matches while pushing them apart for different action classes. Our TDSM significantly outperforms very recent state-of-the-art methods with significantly large margins of 2.36%-point to 13.05%-point, demonstrating superior accuracy and scalability in zero-shot settings through effective skeleton-text matching.
@@ 1 Introduction
Figure 1: Overview of our Triplet Diffusion for Skeleton-Text Matching (TDSM) pipeline versus previous methods. While the previous methods rely on direct alignment between skeleton and text latent spaces, thus suffering from modality gaps that limit generalization, our TDSM pipeline overcomes this challenge by utilizing the cross-modality alignment power of diffusion models, establishing a more unified and robust skeleton-text representation for effective cross-modal matching.
@@
Despite these benefits, the fully supervised skeleton-based action recognition methods [47, 7, 11, 52, 10, 50, 5, 8] tend to perform well, but annotating every possible action is impractical for a large number of possible action classes. In addition, retraining models for new classes incurs a significant cost. So, zero-shot skeleton-based action recognition (ZSAR) [37, 14, 51, 53, 26, 6, 46, 23, 25] addresses this issue by enabling predictions for unseen actions without requiring explicit training data, making it valuable for applications such as surveillance, robotics, and human-computer interaction, where continuous learning is infeasible [44, 13]. Importantly, ZSAR is possible because human actions often share common skeletal movement patterns across related actions. By leveraging these shared patterns, ZSAR methods align pre-learned skeleton features with text-based action descriptions, allowing the models to extrapolate from seen actions to unseen ones. This alignment-based approach reinforces the model’s discriminative power, ensuring scalability and reliable zero-shot recognition in real-world scenarios. However, achieving the effective alignment between skeleton data and text features entails significant challenges. While skeleton data captures temporal and spatial motion patterns, the text descriptions for action labels carry high-level semantic information. This modality gap makes it difficult to align their corresponding latent spaces effectively, thus hindering the generalization learning for unseen actions.
Diffusion models [35, 12] have demonstrated strong cross-modal alignment capabilities by incorporating conditioning signals such as text, images, audio, or video to guide the generative process. This conditioning mechanism enables precise cross-modality alignment, ensuring that generated outputs adhere closely to the given condition. Inspired by this property, we propose a novel framework: Triplet Diffusion for Skeleton-Prompt Matching (TDSM), which firstly adopts diffusion models to ZSAR by conditioning the denoising process on text prompts. Our approach utilizes the reverse diffusion process to implicitly align skeleton and text features within a shared latent space, overcoming the challenges of direct feature space alignment.
@@
Fig. 1 illustrates the key differences between previous methods and our proposed method. The previous methods [37, 14, 51, 53, 26, 6, 46, 23, 25] attempt to directly align skeleton and text features within separate latent spaces. However, this approach struggles with generalization due to the inherent modality gap between skeletal motion and textual semantics. On the other hand, our TDSM leverages a reverse diffusion training scheme to implicitly align skeleton features with their corresponding text prompts, producing discriminatively fused representations within a unified latent space. More specifically, our TDSM learns to denoise noisy skeleton features conditioned on the corresponding text prompts, embedding the prompts into a unified skeleton-text latent space to better capture the semantic meaning of action labels. This implicit alignment mitigates the limitations of direct latent space mapping while enhancing robustness. Additionally, we introduce a triplet diffusion (TD) loss, which encourages tighter alignment for correct skeleton-text pairs and pushes apart incorrect ones, further improving the model’s discriminative power. As an additional benefit, the stochastic nature of diffusion process, driven by the random noise added during training, acts as a natural regularization mechanism. This prevents overfitting and enhances the model’s ability to generalize effectively to unseen actions. Our contributions are threefold:
@@
  • We firstly present a diffusion-based action recognition with zero-shot learning for skeleton inputs, called a Triplet Diffusion for Skeleton-Text Matching (TDSM) which is the first framework to apply diffusion models and to implicitly align the skeleton features with text prompts (action labels) by fully taking the advantage of excellent text-image correspondence learning in generative diffusion process, thus being able to learn fused discriminative features in a unified latent space;
@@
  • We introduce a reformulated triplet diffusion (TD) loss to enhance the model’s discriminative power by ensuring accurate denoising for correct skeleton-text pairs while suppressing it for incorrect pairs;
@@ 2.1 Zero-shot Skeleton-based Action Recognition
Zero-shot Skeleton-based Action Recognition (ZSAR) aims to recognize human actions from skeleton sequences without requiring labeled training data for unseen action categories. Most of the existing works focus on aligning the skeleton latent space with the text latent space. These approaches can be categorized broadly into VAE-based methods [37, 14, 25, 26] and contrastive learning-based methods [51, 53, 6, 23, 46].
@@
While most existing methods rely on direct alignment between skeleton and text latent spaces, they often struggle with generalization due to inherent differences between the two modalities. In contrast, our TDSM leverages diffusion models for alignment rather than generation. By conditioning the reverse diffusion process on action labels, we guide the denoising of skeleton features to implicitly align them with their corresponding semantic contexts of action labels. This enables more robust skeleton-text matching and improves generalization to unseen actions.
@@ 2.2 Diffusion Models
Diffusion models have become a fundamental to generative tasks by learning to reverse a noise-adding process for original data recovery. Denoising Diffusion Probabilistic Models (DDPMs) [16] introduced a step-by-step denoising framework, enabling the modeling of complex data distributions and establishing the foundation for diffusion-based generative models. Building on this, Latent Diffusion Models (LDMs) [35, 12] improve computational efficiency by operating in lower-dimensional latent spaces while maintaining high-quality outputs. LDMs have been successful in various generation tasks (e.g., text-to-image, text-to-video, image-to-video), showing the potential of diffusion models for cross-modal alignment tasks. In the reverse diffusion process, LDMs employ a denoising U-Net [36] where text prompts are integrated with image features through cross-attention blocks, effectively guiding the model to align the two modalities. Further extending this line of research, Diffusion Transformers (DiTs) [32] integrate transformer architectures into diffusion processes. In this work, we leverage the aligned fusion capabilities of diffusion models, focusing on the learning process during reverse diffusion rather than their generative power. Specifically, we utilize a DiT-based network as a denoising model, where text prompts guide the denoising of noisy skeleton features. This approach embeds text prompts into the unified latent space in the reverse diffusion process, ensuring robust fusion of the two modalities and enabling effective generalization to unseen actions in zero-shot recognition settings.
@@
3 Preliminaries: Diffusion Process
Diffusion models [16, 35, 32] are a class of generative models that progressively denoise a noisy sample to generate a target distribution. The diffusion process [16] consists of two main stages: a forward process and a reverse process.
Forward diffusion process. In the forward process, Gaussian noise is incrementally added to the target data $\mathbf{x}_{0}$ over discrete timesteps $t$, gradually transforming it into a Gaussian noise distribution $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. This can be formulated as: $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$, where $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ follows a Markov chain that progressively corrupts $\mathbf{x}_{0}$ into noise, and $\beta_{t}$ controls the noise schedule. By reparameterizing the forward process, $\mathbf{x}_{t}$ can be directly expressed in terms of $\mathbf{x}_{0}$ as: $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ controls the noise level at step $t$.
Reverse diffusion process. The reverse process learns to recover the original data $\mathbf{x}_{0}$ from a noisy sample by estimating the denoising step conditioned on previous timesteps as: $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\bm{\mu}_{\theta}(\mathbf{x}_{t},t),\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t},t))$, where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ follows a Gaussian conditional distribution that gradually reconstructs $\mathbf{x}_{0}$ by estimating the denoised mean $\bm{\mu}_{\theta}(\mathbf{x}_{t},t)$ and covariance $\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t},t)$. Here, $\bm{\mu}_{\theta}$ and $\mathbf{\Sigma}_{\theta}$ are predicted by a neural network trained to approximate the inverse of the forward noise addition, enabling the model to progressively refine $\mathbf{x}_{t}$ into a clean representation.
Objective function. To train the reverse process, the objective function is derived by minimizing the variational bound on the negative log-likelihood $\mathbb{E}\left[-\log p_{\theta}(\mathbf{x}_{0})\right]$. It can be simplified by reparameterizing $\bm{\mu}_{\theta}$ as a noise prediction network $\epsilon_{\theta}$, leading to the following loss as: $\mathcal{L}_{\text{diff}}=\|\epsilon_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}\|_{2}$.
Cross-modality conditioning in diffusion models. An important property of diffusion models is their ability to incorporate conditioning signals $\mathbf{c}$ such as text, images, audio, or video to guide the generative process. This conditioning mechanism enables strong cross-modality alignment, allowing the models to generate outputs that closely align with the given conditions $\mathbf{c}$. To integrate conditioning into the diffusion process, the noise prediction model $\epsilon_{\theta}$ is conditioned on $\mathbf{c}$, modifying the objective function as follows: $\mathcal{L}_{\text{diff}}=\|\epsilon_{\theta}(\mathbf{x}_{t},t;\mathbf{c})-\bm{\epsilon}\|_{2}$. This formulation ensures that the denoising process is guided by the conditions $\mathbf{c}$, enforcing the alignment between the generated output and the conditioning signals $\mathbf{c}$.
@@ 4.1 Overview of TDSM
$\mathcal{D}_{\text{train}}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N},y_{i}\in\mathcal{Y},$ (1)
@@
$\mathcal{D}_{\text{test}}=\{(\mathbf{X}_{j}^{u},y_{j}^{u})\}_{j=1}^{N_{u}},y_{j}^{u}\in\mathcal{Y}_{u},$ (2)
@@
Fig. 2 provides an overview of our training framework of TDSM. As detailed in Sec. 4.2, the pretrained skeleton encoder $\mathcal{E}_{x}$ and text encoder $\mathcal{E}_{d}$ embed the skeleton inputs $\mathbf{X}$ and prompt input $\mathbf{d}$ with an action label $y$ into their respective feature spaces, producing the skeleton feature $\mathbf{z}_{x}$ and two types of text features: the global text feature $\mathbf{z}_{g}$ and the local text feature $\mathbf{z}_{l}$. The skeleton feature $\mathbf{z}_{x}$ undergoes the forward process, where noise $\bm{\epsilon}$ is added to it. In the reverse process, as described in Sec. 4.3, the Diffusion Transformer $\mathcal{T}_{\text{diff}}$ serves as the noise prediction network $\epsilon_{\theta}$, predicting the noise $\hat{\bm{\epsilon}}$. The training objective function that ensures the TDSM to learn robust discriminative power is discussed in Sec. 4.3. Finally, Sec. 4.4 explains the strategy used during the inference phase to predict the correct label ${\hat{y}}^{u}$ for unseen actions.
@@ 4.2 Embedding Skeleton and Prompt Input
Following the LDMs [35, 12], we perform the diffusion process in a compact latent space by projecting both skeleton data and prompt into their respective feature spaces. For skeleton data, we adopt GCNs as the architecture for the skeleton encoder $\mathcal{E}_{x}$ that is trained on $\mathcal{D}_{\text{train}}$ using a cross-entropy loss:
@@
For text encoder $\mathcal{E}_{d}$, we leverage the text prompts to capture rich semantic information about the action labels. Each ground truth (GT) label $y_{p}=y$ is associated with a prompt $\mathbf{d}_{p}$, while a randomly selected wrong label (negative sample) $y_{n}\in\mathcal{Y}\setminus\{y_{p}\}$ is assigned a prompt $\mathbf{d}_{n}$. To encode these prompts, we utilize a pretrained text encoder, such as CLIP [34, 19], which provides two types of output features: a global text feature $\mathbf{z}_{g}$ and a local text feature $\mathbf{z}_{l}$. The text encoder’s output for a given prompt d can be expressed as:
@@ 4.3 Diffusion Process
$\mathbf{z}_{x,t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{x}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon}.$ (6)
@@
$\hat{\bm{\epsilon}}=\mathcal{T}_{\text{diff}}\left(\mathbf{z}_{x,t},t;\mathbf{z}_{g},\mathbf{z}_{l}\right).$ (7)
Using the shared weights in $\mathcal{T}_{\text{diff}}$, we predict $\hat{\bm{\epsilon}}_{p}$ for positive features $\left(\mathbf{z}_{g,p},\mathbf{z}_{l,p}\right)$ and $\hat{\bm{\epsilon}}_{n}$ for negative features $\left(\mathbf{z}_{g,n},\mathbf{z}_{l,n}\right)$. $\mathcal{T}_{\text{diff}}$ builds upon DiT [32], which has been well-validated for cross-modality alignment in image-text tasks. We adopt it to the skeleton-text domain by: (i) reducing the number of blocks/channels to accommodate the relatively small-scale skeleton data, and (ii) incorporating both global and local text embeddings to enhance skeleton-text alignment.
Triplet diffusion (TD) loss. The overall training objective combines a diffusion loss and our reformulated TD loss, which is inspired by the conventional triplet loss [17], to promote both effective noise prediction and discriminative alignment (fusion). The total loss is defined as:
@@
where $\mathcal{L}_{\text{diff}}$ ensures accurate denoising, and $\mathcal{L}_{\text{TD}}$ enhances the ability to differentiate between correct and incorrect label predictions. The diffusion loss $\mathcal{L}_{\text{diff}}$ is given by: